a woman makes 15% more than her husband.Together they make 69,875 per year. what is the husbands annual salary?
Answer by math problem solving(30)

let x be her husband's annual income.
Then she earns x + 15%*x.
As their total earning is $69875, we have
x + x + 15%*x = 69,875
Solving for x, we have
x + x + 0.15x = 69875
2.15x = 69875
x = 69875/2.15
x=32,500
So the husband's annual salary is $32,500.